# LPSim: Large Scale Multi-GPU Parallel Computing based Regional Scale Traffic Simulation Framework

### v1

@@ 4.2 Experimental Results
Thread predication in GPUs deactivates threads not following a branch, impacting performance by limiting active thread count in a warp. Since a warp processes a single instruction collectively, the fraction of active threads directly influences efficiency. The dotted line in figures represents the maximum warp-level performance, with actual performance falling below this threshold. As is shown in figure 13(a), the dots are well below the dotted line indicating a 10× loss in performance due to thread predication. To address such a significant performance loss, we sorted the input OD pairs by departure time. This approach is depicted in the new roofline results shown in Figure 13(b), where the gap between the performance markers and the theoretical maximum (dotted line) noticeably narrows, reducing the performance loss from a factor of 10 to just 2. This improvement is largely due to the sorting making threads within the same warp more likely to execute similar instructions, significantly reducing the inefficiencies associated with thread predication. For simulations of vehicles within the same area, minimizing the number of instructions is preferable as it leads to reduced resource wastage, thereby accelerating the computation process.
@@
Figure 13: The Instruction Roofline for our simulator.
@@ 5 Conclusion
The experimental results, validated against real-world data, demonstrate LPSim’s superior accuracy in replicating traffic dynamics. The framework’s performance analysis reveals significant advantages in using multiple GPUs over a single GPU setup, including scalability and efficiency in handling large-scale traffic simulations.

### v2

@@ 4.2 Experimental Results
Figure 12(a) shows the performance analysis using the Roofline model, where the dots are well below the dotted line representing the maximum warp-level performance. This indicates a 10× loss in performance due to thread predication. To address this significant performance loss, we sorted the input OD pairs by departure time. The new roofline results, shown in Figure 12(b), reveal that the gap between the performance markers and the theoretical maximum (dotted line) noticeably narrows, reducing the performance loss from a factor of 10 to just 2. This improvement is largely due to the sorting making threads within the same warp more likely to execute similar instructions, significantly reducing the inefficiencies associated with thread predication. Minimizing the number of instructions for simulations of vehicles within the same area is preferable as it leads to reduced resource wastage, thereby accelerating the computation process.
@@
Figure 12: The Instruction Roofline for our simulator. The black line represents the theoretical maximum instruction throughput on a warp basis, calculated based on the A100’s architecture, which comprises 108 Streaming Multiprocessors (SMs) with four sub-partitions each, integrating a single warp scheduler capable of dispatching one instruction per cycle within each sub-partition. Consequently, the theoretical maximum instruction throughput on a warp basis is calculated as 108×4×1×1.41 GHz, amounting to 609.12 GIPS. Performance metrics are determined by the methodology of Yang et al. (2019), normalized to warp-level, and rescaled into gigasectors per second (GSECT/s) based on the sector size. The red line indicates the efficiency of memory access patterns. A warp-level load can result in 1 to 32 sector transactions, making the x-axis digit meaningful. A value of 1 indicates “stride-0” access, where all threads in a warp reference a single memory location, generating only one transaction. Conversely, scenarios like random access or striding beyond 32 bytes (“stride-8”) can lead to the maximum 32 transactions. For unit-stride (“stride-1”) access, typical of FP32 or INT32 operations, the global load/store (LD/ST) intensity stands at 1/4. Blue dots represent non-predicated instructions for each level of the memory hierarchy, highlighting the limited data locality and speedup from sorting. Gold dots are non-predicated loads per global memory access, highlighting the loss in near-unit stride access from sorting. The dotted line represents the total (including predicated) instruction throughput. Proximity of blue dots to the highlighted line quantifies the impact of predication.
4) Simulation time with different demand sizes. We selected demand sizes of 3M, 6M, 12M, 20M, and 24M demands randomly from the full-day SFCTA demand as described in Section 4.1 to be tested on AWS P3 instances with various numbers of NVIDIA V100 GPUs. We can see that for all the demand pairs, we gained time benefits from increasing GPU numbers from 1 to 2. And for demand of 12M and 20M, we can see time benefit from using 1 GPU all the way to 4 GPUs. And for 24M demand, it can only be simulated with more than 4 GPUs.
Figure 13: Simulation Time for Different Demands with different number of GPUs
Table 7 demonstrates a reduction in simulation time as the setup scales from 1 to 4 GPUs in both balanced and the unbalanced partition with demands of more than 12M, highlighting an enhancement in multi-GPU simulation performance. However, the increased runtime observed when transitioning from 4 to 8 GPUs in the balanced and unbalanced scenario suggests that communication becomes a bottleneck in scenarios where the number of GPUs increases without a corresponding enlargement of the problem size.
Table 7: Simulation Times for Different Partitions
Demand Size Partition Type Simulation Time with Different Number of GPUs
1 GPU 2 GPUs 4 GPUs 8 GPUs
3,000,000 Balanced Partition 381997 353399 439829 689898
Unbalanced Partition 381997 343435 446019 678993
6,000,000 Balanced Partition 752695 610608 653128 928397
Unbalanced Partition 752695 597259 632208 868127
12,000,000 Balanced Partition 1535082 1182821 1087738 1306006
Unbalanced Partition 1535082 1134121 1111049 1304821
20,000,000 Balanced Partition 2554437 1746744 1459479 1639141
Unbalanced Partition 2554437 1768196 1717089 1635972
24,000,000 Balanced Partition n/a n/a 1744777 1938588
Unbalanced Partition n/a n/a 2011620 1928708
@@ 5 Conclusion
The experimental results, validated against real-world data, demonstrate LPSim’s accuracy Jiang et al. [2024] in replicating traffic dynamics. The framework’s performance analysis reveals significant advantages in using multiple GPUs over a single GPU setup, including scalability and efficiency in handling large-scale traffic simulations.